A man is holding an 8.00-kg vacuum cleaner at arm'
Answer #1

Given that 3= 90-38 = 600 830.550 m, m = 8.00 kg, g=9.81 m/s The torque (5) = ngr sine =(8.00X9.91)(0.550) since Answer:- 37.